Interesting ques..

Interesting ques..

A circular conductin loop of radius r0 &having resistance per unit length  is placed vertically in  magnetic field B which is constant in space &time.The ends of the loop are crossed and pulled in opposite directions with velo v such dat the loop always remain circular and the radius of the loop goes on decreasing..then

1)radius of loop changes with  time t as???

2)Find the current induced in the loop..


Grade:10

2 Answers

ROSHAN MUJEEB
3 years ago
Perimeter is decreasing at a rate of 2v
ddt(2πr)=2v⇒drdt=vπ
∴r=(r0−vπt)
ϕ=Bπr2⇒ɛ=∣∣∣−dϕdt∣∣∣=B2πrdrdt
∴ɛ=2Bπ(r0−vπt)vπ=2Bv(r0−vπt)
I=ϕR=2Bvrλ.2πr=Bvπλ
ROSHAN MUJEEB
3 years ago
Perimeter is decreasing at a rate of 2v
ddt(2πr)=2v⇒drdt=vπddt(2πr)=2v⇒drdt=vπ
∴r=(r0−vπt)∴r=(r0-vπt)
ϕ=Bπr2⇒ɛ=∣∣∣−dϕdt∣∣∣=B2πrdrdtϕ=Bπr2⇒ɛ=|-dϕdt|=B2πrdrdt
∴ɛ=2Bπ(r0−vπt)vπ=2Bv(r0−vπt)∴ɛ=2Bπ(r0-vπt)vπ=2Bv(r0-vπt)
I=ϕR=2Bvrλ.2πr=Bvπλ